Points P and Q are between point A and B with P closer to A. If A(-1,2) and B(8,20) find the coordinates of P and Q such that AP:PQ:QB=2:3:4.
Answer by Alan3354(69443)

Points P and Q are between point A and B with P closer to A. If A(-1,2) and B(8,20) find the coordinates of P and Q such that AP:PQ:QB=2:3:4.
--------------
2+3+4 = 9
P is 2/9 of the distance from A to B
Q is (2+3)/9 = 5/9 of the distance from A to B
---------------
Do x and y separately.
For x: AB distance = 9
P = -1 + (2/9)*9 = 1
Q = -1 + (5/9)*9 = 4
--------------------
For y: AB distance = 18
P = 2 + (2/9)*18 = 6
Q = 2 + (5/9)*18 = 12
===========================
P(1,6)
Q(4,12)